# Does Synthetic Data Help Named Entity Recognition for Low-Resource Languages?

### v2

@@ Abstract
In this paper, we explore whether synthetic datasets generated by large language models are useful for low-resource named entity recognition, considering 11 languages from diverse language families. Our results suggest that synthetic data created with seed human labeled data is a reasonable choice when there is no available labeled data, and is better than using automatically labeled data. However, a small amount of high-quality data, coupled with cross-lingual transfer from a related language, always offers better performance.
@@ 1 Introduction
Named Entity Recognition (NER) for low-resource languages aims to produce robust systems for languages with limited labeled training data available, and has been an area of increasing interest within natural language processing (NLP). Two common approaches to address this data scarcity are cross-lingual transfer and data augmentation/synthesis; recent research has in particular explored the usefulness of large language models (LLMs) for such data augmentation and synthetic data creation in NLP Whitehouse et al. (2023); Li et al. (2023), while their use for NER is also emerging Bogdanov et al. (2024).
In this background, we propose LLM-based synthetic data generation using a small amount of gold examples (Figure 1) as an alternative to relying on automatically created datasets for low-resource NER. With experiments covering 11 languages, we show that
@@
    Synthetic data generated by prompting an LLM with a few high quality examples (Figure 1) could be better than using automatically labeled datasets when training low-resource NER models.
Figure 1: High-level overview of our data generation process. We use multilingual large language models to generate new NER datapoints on the basis of a handful of high quality data points. See Section 3.1 for more.
@@ 2 Related Work
NER in low resource settings has long been a topic of interest in NLP. Significant research examines cross-lingual transfer from a high resource source language to a lower-resource target language for the task Rahimi et al. (2019); Mueller et al. (2020); Zeng et al. (2022); Zhao et al. (2022); Yang et al. (2022); Zhou et al. (2022), while other approaches have explored the creation of synthetic datasets through e.g. parallel corpora or machine translation Mayhew et al. (2017); Ni et al. (2017); Pan et al. (2017); Xie et al. (2018); Liu et al. (2021); Yang et al. (2022); Fetahu et al. (2022).
More recent work has explored using LLMs such as GPT-3.5 and GPT-4 as data generators for NER Bogdanov et al. (2024); Heng et al. (2024). We build on such work, but differ from their methods. Our data generation process uses high quality examples as seeds, and we not only evaluate different LLMs (both open and closed-source), but also experiment with 11 languages covering three language families and five base scripts.
@@ 3.1 Synthetic Data Generation:
  • We submit this prompt as input to the LLM, and extract any correctly-formatted data points from its response;
@@
We compare three LLMs as our source of synthetic data: GPT-4111We use gpt-4-turbo, and all data generation with the model was conducted between September and December 2024. Achiam et al. (2023), which we assume to be the state of the art; Llama-3.1-8B-Instruct Dubey et al. (2024), as a much smaller, open-source instruction-tuned model; and finally, aya-expanse-32b Dang et al. (2024), as a larger open source multilingual LLM.
3.2 Training NER models:
  1. 1.
    In the first setting, we use our data to train an NER model from scratch, by fine-tuning the pre-trained XLM-RoBERTa-large on target language NER data.
  2. 2.
    In the second setting, we first fine-tune the model on the high quality NER data in a language related source language [^2] See Table 2 in Appendix B for the full list of chosen related languages for all the target languages., and then further fine-tune this NER model on our synthetic or organic target language data.
Languages & Datasets:
We focus on 11 languages from diverse language families: Tamil, Kannada, Malayalam, Telugu (Dravidian), Kinyarwanda, Swahili, Igbo, Yoruba (Niger-Congo), Swedish, Danish and Slovak (Indo-European). Of these, Igbo, Yoruba, and Kinyarwanda are not among the 100 languages in the XLM-Roberta pre-training corpus. We use the Universal NER dataset Mayhew et al. (2024) as our high quality, manually annotated dataset for Swedish, Danish and Slovak; MasakhaNER2 Adelani et al. (2022) for Kinyarwanda, Swahili, Igbo and Yoruba; and the Naamapadam dataset Mhaske et al. (2023) for Tamil, Kannada, Malayalam and Telugu.
While the first two datasets are completely manually annotated, the train and validation splits of the Naamapadam dataset are constructed using parallel corpora, and thus contain some noise. Nevertheless, we choose it as our organic dataset, as (i) its test sets, which contain 500-1000 datapoints per language, are completely manually annotated, and (ii) it remains the largest NER resource for these four languages. Crucially, all of these datasets cover largely identical NER categories, allowing for comparisons between them.
Additionally, we compare models trained on LLM-generated data with those trained using WikiANN Pan et al. (2017); Rahimi et al. (2019), a large, automatically created NER dataset based on Wikipedia cross-linking, as it covers the 11 languages we study. This dataset represents a different form of synthetic data—one generated not from LLMs, but instead from scraping knowledge bases. Although the dataset has no manual annotations, it is frequently used as a standard low-resource NER benchmark Schmidt et al. (2022); Asai et al. (2024).
4 Results
4.1 Synthetic Data Generation
We generate the synthetic datasets following the process described in Section 3.1. While model responses from GPT-4 are almost always usable, we found recurring errors in responses from the other two models. Some of these errors are described in Table 1 in Appendix A; we discard such instances when compiling our synthetic datasets from model responses. The average percentage of usable training datapoints from GPT-4, Llama-3.1 and aya-expanse are 97%, 59.3% and 11.7% respectively. [^3] Llama-3.1’s rejected datapoints are often incomplete due to hitting new token limits, suggesting potentially higher capabilities under higher token limits. We assess the overall quality and viability of this synthetic data by measuring the performance of an NER model on a high quality, manually-annotated test set, when trained on the synthetic data.
4.2 Training on Synthetic Data
Figure 2 shows our results when using synthetic data from different models, in both the ner from scratch and ner fine-tuning settings. While the models trained on organic data in the ner from scratch setting always perform better than synthetic data based models, we find that models trained on GPT-4-generated data come the closest to models trained on organic data. We also find that more synthetic data is not necessarily useful; for some languages, we see a saturation after about 1000 data points, and for some, we also notice a drop in performance with more data.
Perhaps more surprisingly, in the ner fine-tuning setting, we notice that zero-shot transfer from a related language outperforms the same models after they have been further fine-tuned on synthetic target language data. This suggests that in some cases where an NER model for a related language exists, synthetic data in target languages may actually be detrimental to overall performance.
(a) ner from scratch Setting
(b) ner fine-tuning Setting
Figure 2: NER model performance when trained on increasingly large subsets of training data. aya-expanse-32b and Llama-3.1-8B-Instruct produced lower amounts of usable data; this is why they do not extend as far as organic or GPT-4-produced data in fine-tuning data size. Performance at Fine-tuning Dataset Size = 0, only present in the ner fine-tuning setting, indicates zero-shot performance of a related-language NER model.
Comparison with WikiAnn:
In most cases, when dataset size is comparable, training on WikiANN data in the ner from scratch setting yielded NER models that perform considerably worse than those trained on synthetic data from GPT-4. For the four Niger-Congo languages, GPT-4 generated data gave superior results even in the ner fine-tuning setting (see Table 3 in Appendix C for the detailed results).
5 Conclusions and Discussion
  1. 1.
  2. 2.
    In many cases, zero-shot transfer from a related-language NER model is a high baseline, and that further training such a model on synthetic data may even lower the performance.
  3. 3.
    Despite the fact that it falls short of manually annotated data, LLM-generated data often still yields better model performance than WikiANN, especially for the more low-resource languages among the ones we studied. This echoes the findings by Lignos et al. (2022), who arrive at similarly negative findings around the data quality of WikiANN.
6 Limitations
Although we experimented with many languages, the nature of the NER datasets used is relatively simple, containing only three or four entity categories (persons, locations, organizations and dates). Thus, we don’t know if the general conclusions, especially about the quality of synthetic data, will extend to scenarios where there are many entity categories. While we did study datasets covering more than one language family, the selection of language is far from extensive, and is also constrained by the availability of human labeled test data. Finally, to keep the experiments under control, we explored a limited set of methods for fine-tuning and synthetic data generation. Our findings should be viewed after taking these aspects into consideration.
7 Ethics and Broader Impact
We used publicly available datasets with human-annotated and automatically labeled data, and also created synthetically generated datasets as a part of this work. The models built using such artificially created datasets should always be validated with a human-labeled data. We did not involve any human participants in this study. All the code and generated datasets is provided at this github repository to support reproducible research: https://anonymous.4open.science/r/low-resource-syn-ner-A1C7/.
Appendix A Synthetic Data Generation
@@
Table 1: Examples of different types of responses from the synthetic data-generating LLMs tested.
@@
For GPT-4, we used the OpenAI API’s functionalities for structured outputs to ensure that outputs were formatted as JSON strings. For the open-sourced models, we experimented with using transformers-compatible libraries for obtaining structured outputs from LLMs, but ultimately found better results simply specifying the JSON requirement in the model and system prompt. For the open-sourced models, we used the following system prompt:
@@
For GPT-4, we used the following (minimally different) system prompt, on the assumption that specifying output mode in the system prompt was less important on account of the API’s structured output functionalities:
@@
We ran all open-sourced models using vLLM Kwon et al. (2023), with a temperature setting of 0.8, maximum new token limit of 4096 new tokens, and nucleus sampling value of 0.8. Calls to GPT-4 were made using default hyperparameters.
@@ Appendix B Related-Language Model Details
Table 2: List of related languages used in the ner fine-tuning setting for each target language. *English is not closely related to Slovak, but given the absence of another highly related language among the 11 target languages, it was chosen as the language for the base NER model to be fine-tuned.
@@ Appendix C Full Results of WikiANN Comparison
The WikiANN dataset is a massively multilingual NER benchmark, comprising data from 176 languages Pan et al. (2017); Rahimi et al. (2019). [^4] As Lignos et al. (2022) also note, strictly speaking, the original version of WikiANN put together by Pan et al. (2017) contains data from 282 languages; the version of the dataset commonly downloaded from Huggingface, however, and put together by Rahimi et al. (2019), contains data from 176 languages. In this work, we refer to the latter when referring to the WikiANN dataset. Table 3 shows the full list of comparisons between NER model performance when trained on organic data, GPT-4-produced data, and WikiANN data. The sizes of the WikiANN train sets vary significantly between different languages, meaning we often cannot assess the quality of the data in the context of training sets containing over 1000 datapoints (e.g. Kannada and Yoruba, whose WikiANN train sets contain only 100 datapoints). In such cases, however, we compare model performance when trained on equally small amounts of organic or LLM-produced synthetic data.
@@
GPT-4 0.29 0.29 4889
Universal NER 0.80 0.82 5000
Table 3: Performance of NER models trained on WikiANN, synthetic data from GPT-4, and high quality ‘organic’ data, for all 11 languages. n.f.s: ner from scratch setting; n.f.t: ner fine-tuning setting.

### v3

@@ Abstract
We explore whether synthetic datasets generated by large language models using a few high quality seed samples are useful for low-resource named entity recognition, considering 11 languages from three language families. Our results suggest that synthetic data created with such seed data is a reasonable choice when there is no available labeled data, and is better than using entirely automatically labeled data. However, a small amount of high-quality data, coupled with cross-lingual transfer from a related language, always offers better performance. [^1] Data and code available at: https://github.com/grvkamath/low-resource-syn-ner.
@@ 1 Introduction
Named Entity Recognition (NER) for low-resource languages aims to produce robust systems for languages with limited labeled training data available, and has been an area of increasing interest within natural language processing (NLP) over the past decade. Two common approaches to address this data scarcity are cross-lingual transfer and data augmentation/synthesis; recent research has in particular explored the usefulness of large language models (LLMs) for such data augmentation and synthetic data creation in NLP Whitehouse et al. (2023); Li et al. (2023), while their use for NER is also emerging Bogdanov et al. (2024); Dao et al. (2025).
Figure 1: High-level overview of our data generation process. We use multilingual large language models to generate new NER data points on the basis of a handful of high quality human labeled data points. See Section 3.1 for more.
In this background, we propose LLM-based synthetic data generation using a small amount of gold examples (Figure 1) as an alternative to relying on automatically created datasets for low-resource NER. With experiments covering 11 languages from 3 language families—Danish, Swedish and Slovak from the Indo-European language family; Swahili, Kinyarwanda, Yoruba and Igbo from the Niger-Congo language family, and Kannada, Malayalam, Tamil, and Telugu from the Dravidian language family, we show that:
@@
    Synthetic data generated by prompting an LLM with a few high quality (generally human labeled) examples (Figure 1) could be better than using automatically labeled datasets when training low-resource NER models.
@@ 2 Related Work
NER in low resource settings has long been a topic of interest in NLP. Significant research examines cross-lingual transfer from a high resource source language to a lower-resource target language for the task Rahimi et al. (2019); Mueller et al. (2020); Zeng et al. (2022); Zhao et al. (2022); Yang et al. (2022); Zhou et al. (2022), while other approaches have explored the creation of synthetic datasets through e.g. parallel corpora or machine translation Mayhew et al. (2017); Ni et al. (2017); Pan et al. (2017); Xie et al. (2018); Liu et al. (2021); Yang et al. (2022); Fetahu et al. (2022). There are also large existing automatically constructed multilingual NER datasets that rely on sources such as Wikipedia Pan et al. (2017); Krishnan et al. (2021); Malmasi et al. (2022), some of which have become a part of large multilingual benchmarks Asai et al. (2024).
More recent work has explored using LLMs as data generators for NER Bogdanov et al. (2024); Heng et al. (2024); Evuru et al. (2024). We build on such work, but differ from their methods. Our data generation process uses high quality, human validated examples as seeds, and we not only evaluate different LLMs (both open and closed-source) as synthetic data generators, but also experiment with 11 languages covering three language families and five base scripts. To our knowledge, this is the first attempt to explore using large language models for synthetic data generation in low-resource NER, and the first to cover >10>10 languages.
@@ 3.1 Synthetic Data Generation:
  • We submit this prompt as input to the LLM, and extract the correctly-formatted data points from its response;
@@
Table 1: Examples of different types of responses from the synthetic data-generating LLMs tested, across languages
We compare three LLMs as our source of synthetic data: GPT-4.1222We use gpt-4.1-2025-04-14. Note that in an earlier draft of this work, we used gpt-4-turbo Achiam et al. (2023), when it represented the state-of-the-art; surprisingly, gpt-4-turbo yielded slightly better results. Nevertheless, here we report results on GPT-4.1, to better represent currently available models. OpenAI (2025), which we assume to be the state of the art; Llama-3.1-8B-Instruct Dubey et al. (2024), as a much smaller, open-source instruction-tuned model; and finally, aya-expanse-32b Dang et al. (2024), as a larger open source multilingual LLM.
3.2 Training NER models:
  1. 1.
    In the first setting, we use our synthetic data to train an NER model from scratch, by fine-tuning XLM-RoBERTa-large on target language NER data.
  2. 2.
    In the second setting, we first fine-tune the model on the high quality NER data in a related source language [^3] See Table 2 in Appendix B for the full list of chosen related languages for all the target languages., and then further fine-tune this NER model using our synthetic or organic target language data.
Languages & Datasets:
We focus on 11 languages from three distinct language families: Tamil, Kannada, Malayalam, Telugu (Dravidian), Kinyarwanda, Swahili, Igbo, Yoruba (Niger-Congo), Swedish, Danish and Slovak (Indo-European). Of these, Igbo, Yoruba, and Kinyarwanda are not among the 100 languages in the XLM-Roberta pre-training corpus. We use the Universal NER dataset Mayhew et al. (2024) as our high quality, manually annotated dataset for Swedish, Danish and Slovak; MasakhaNER2 Adelani et al. (2022) for Kinyarwanda, Swahili, Igbo and Yoruba; and the Naamapadam dataset Mhaske et al. (2023) for Tamil, Kannada, Malayalam and Telugu.
While the first two datasets are completely manually annotated, the train and validation splits of the Naamapadam dataset are constructed using parallel corpora, and thus contain some noise. Nevertheless, we choose it as our organic dataset, as (i) its test sets, which contain 500-1000 datapoints per language, are completely manually annotated, and (ii) it remains the largest NER resource for these four languages. Crucially, all of these datasets cover largely identical NER categories, allowing for comparisons between them. The Universal NER and Naamapadam datasets cover persons, locations and organizations as categories; the MasakhaNER2 data covers these three categories, as well as dates.
Additionally, we compare models trained entirely on LLM-generated data with those trained using WikiANN Pan et al. (2017); Rahimi et al. (2019), a large, automatically created NER dataset based on Wikipedia cross-linking, as it covers the 11 languages we study. This dataset represents a different form of synthetic data—one generated not from LLMs, but instead from scraping knowledge bases without any seed data. Although the dataset has no manual annotations, it is frequently used as a standard low-resource NER benchmark Schmidt et al. (2022); Asai et al. (2024).
4 Results
(a) ner from scratch Setting
(b) ner fine-tuning Setting
Figure 2: NER model performance when trained on increasingly large subsets of training data. aya-expanse-32b and Llama-3.1-8B-Instruct produced lower amounts of usable data; this is why they do often not extend as far as organic or GPT-4.1-produced data in fine-tuning data size. In the ner fine-tuning setting, performance at Fine-tuning Training Data Size = 0 indicates zero-shot performance of a related-language NER model.
4.1 Synthetic Data Generation
We generate the synthetic datasets following the process described in Section 3.1. While model responses from GPT-4.1 are generally usable, we found more recurring errors in responses from the other two models. Some of these errors are described in Table 1; we discard such instances when compiling our synthetic datasets from model responses. The average percentage of usable training datapoints from GPT-4.1, Llama-3.1 and aya-expanse are 82.6%, 59.7% and 11.1% respectively. We assess the overall quality and viability of this synthetic data by measuring the performance of an NER model on a high quality, manually-annotated test set, when trained on the synthetic data.
4.2 Training on Synthetic Data
Figure 2 shows our results when using synthetic data from different models, in both the ner from scratch and ner fine-tuning settings. While the models trained on organic data in the ner from scratch setting always perform better than synthetic data based models, we find that the models trained on GPT-4.1-generated data often come the closest to models trained on organic data compared to the other synthetic data sources. Best results with synthetic data based training are seen for Danish, followed by Swahili and Swedish. We also find that more synthetic data is not necessarily useful; for most languages, we see relative saturation after 1000 data points, and in the case of Kannada, we actually notice a drop in performance with more data.
Perhaps more surprisingly, in the ner fine-tuning setting, we notice that zero-shot transfer from a related language usually outperforms the same models after they have been further fine-tuned on synthetic target language data. Fine-tuning the related language NER model with organic data from the target language helped for only Kinyarwanda and Yoruba. This suggests that in some cases where an NER model for a related language exists, synthetic data in target languages may actually be detrimental to overall performance. Models trained on synthetic data from GPT-4.1 do better than those trained on synthetic data from Llama-3.1-8B-Instruct only about half of the time; on the other hand, there are often too few usable aya-expanse-32b datapoints for a fair comparison.
Comparison with WikiAnn:
In addition to comparing models fine-tuned on synthetic data versus organic data, we also looked into the question of whether our synthetic data generation approach offers any benefits over automatically labeled datasets, taking WikiAnn as an example. Training models on WikiANN data leads to higher performance than training on GPT-4.1-generated data only for the four Dravidian languages in our data, but generally leads to significantly worse performance than training on synthetic data from GPT-4.1 for the remaining languages (see Table 3 in Appendix C for detailed results). This holds in both the ner from scratch and ner fine-tuning settings, when data size is comparable; and, in the case of Danish and Swedish, training on WikiANN leads to worse performance even when there is several times more WikiANN data than GPT-4.1-generated data. Overall, we can conclude that using synthetic data following our approach appears to be better than relying on WikiAnn for most languages. This echoes the findings by Lignos et al. (2022), who arrive at similarly negative findings around the data quality of WikiANN, and calls for not considering results on WikiANN as a benchmark for multilingual NER comparisons in the future.
5 Conclusions and Discussion
  1. 1.
  2. 2.
    In many cases, zero-shot transfer from a related-language NER model is a high baseline, and that further training such a model on synthetic data may even lower the performance. We find this to be true in the case of all languages tested except the Yoruba and Swahili. For these two languages, it is worth noting that the overall baselines are lower, presumably because these languages are all lower resource than the others tested. This may explain why synthetic data yields performance gains over the zero-shot baseline, though it does not change the trend of a small amount of manually annotated data yielding far better performance.
  3. 3.
    Despite the fact that it falls short of manually annotated data, LLM-generated data often still yields better model performance than WikiANN, which is automatically extracted from Wikipedia texts.
6 Limitations
Although we experimented with many languages, the nature of the NER datasets used is relatively simple, containing only three or four entity categories (persons, locations, organizations and dates). Thus, we don’t know if the general conclusions, especially about the quality of synthetic data, will extend to scenarios where there are many entity categories. While we did study datasets covering more than one language family, the selection of language is far from extensive, and is also constrained by the availability of human labeled test data. The observations need not necessarily hold true across all language families, naturally. Finally, to keep the experiments under control, we explored a limited set of methods for fine-tuning and synthetic data generation. Our findings should be viewed after taking these aspects into consideration.
7 Ethics and Broader Impact
We used publicly available datasets with human-annotated and automatically labeled data, and also created synthetically generated datasets as a part of this work. The models built using such artificially created datasets should always be validated with a human-labeled data. We did not involve any human participants in this study. All the code and generated datasets is provided at this GitHub repository to support reproducible research: https://github.com/grvkamath/low-resource-syn-ner.
Acknowledgments
This research was conducted at the National Research Council of Canada, thereby establishing a copyright belonging to the Crown in Right of Canada, that is, to the Government of Canada. Gaurav Kamath is supported by a Doctoral Training Award from the Fonds de Recherche du Québec—Société et Culture.
Appendix A Synthetic Data Generation
@@
We prompt GPT-4.1 using OpenAI’s batch API functionality [^4] https://platform.openai.com/docs/guides/batch; for the open-source models, we use the vLLM library Kwon et al. (2023) to run inference.
For GPT-4.1, we used the OpenAI API’s functionalities for structured outputs to ensure that outputs were formatted as JSON strings. For the open-sourced models, we experimented with using transformers-compatible libraries for obtaining structured outputs from LLMs, but ultimately found better results simply specifying the JSON requirement in the model and system prompt. For the open-sourced models, we used the following system prompt:
@@
For GPT-4.1, we used the following (minimally different) system prompt, on the assumption that specifying output mode in the system prompt was less important on account of the API’s structured output functionalities:
@@
We ran both open-sourced models with a temperature setting of 0.8, and nucleus sampling value of 0.8. We initially used a maximum new token limit of 4096 for both models. However, noticing that some of Llama-3.1-8B-Instruct’s unusable datapoints were specifically due to hitting new token limits, we regenerated data from this model with a maximum new token limit of 8192. Calls to GPT-4.1 were made using default hyperparameters.
@@ Appendix B Related-Language Model Details
Table 2: List of related languages used in the ner fine-tuning setting for each target language. *English is not closely related to Slovak, but given the absence of another closely related language among the 11 target languages, it was chosen as the language for the base NER model to be fine-tuned.
@@ Appendix C Full Results of WikiANN Comparison
The WikiANN dataset is a massively multilingual NER benchmark, comprising data from 176 languages Pan et al. (2017); Rahimi et al. (2019). [^5] As Lignos et al. (2022) also note, strictly speaking, the original version of WikiANN put together by Pan et al. (2017) contains data from 282 languages; the version of the dataset commonly downloaded from Huggingface, however, and put together by Rahimi et al. (2019), contains data from 176 languages. In this work, we refer to the latter when referring to the WikiANN dataset. Table 3 shows the full list of comparisons between NER model performance when trained on organic data, GPT-4.1-produced data, and WikiANN data. The sizes of the WikiANN train sets vary significantly between different languages, meaning we often cannot assess the quality of the data in the context of training sets containing over 1000 datapoints (e.g. Kannada and Yoruba, whose WikiANN train sets contain only 100 datapoints). In such cases, however, we compare model performance when trained on equally small amounts of organic or LLM-produced synthetic data.
@@
GPT-4.1 0.29 0.29 4508
Universal NER 0.77 0.80 5000
Table 3: Performance of NER models trained on WikiANN, synthetic data from GPT-4.1, and high quality ‘organic’ data, for all 11 languages. n.f.s: ner from scratch setting; n.f.t: ner fine-tuning setting.